how long is the hypotenuse of a 30-60-90 triangle .if the shorter leg measures 10 cm ?
Answer by oberobic(2304)

The ratio of the sides of a 30-60-90 triangle are known to be 1:2:sqrt(3).
The shorter leg is 10 cm.
Therefore, the hypotenuse is 2*10 = 20 cm.
.
Check the solution with the Pythagorean Theorem.
c^2 = a^2 + b^2
20^2 = 10^2 + (10*sqrt(3))^2
400 = 400
True.
.
The hypotenuse of the triangle is 20 cm.
.
Done.